theres a right triangle with 2 equal sides the area is 200cm squared what is the area of the missing side
Answer by sudhanshu_kmr(1152)


1/2 * x*x = 200
=> x= 20
h = root(20^2 + 20^2)
= root(800)
= 20root(2)
so, sides are 20,20 and 20(root 2).